# Hölder regularity of harmonic functions on metric measure spaces

### v2

@@ 4 Poisson equation on unbounded MMD spaces
$\leq\sum_{n=N}^{\infty}\int_{B(x,2^{n+1}\Psi^{-1}(t))\backslash B(x,2^{n}\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y),t)\right)m(\mathrm{d}y)$
@@
Recall that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent if for any non-negative function $f\in L^{1}(X;m)$, we have either $Gf=+\infty$ or $Gf=0$ $m$-a.e., where $G$ is the Green operator, see [26, Equation (1.6.2)]. In our setting, the Dirichlet form is recurrent as follows.
@@
We use some techniques from resistance forms. The theory of resistance forms was developed by Kigami [44]. To begin with, let us introduce effective resistances. Let $A,B$ be two non-empty subsets of $X$. The effective resistance between $A$ and $B$ is defined as follows.
@@
In the strongly recurrent setting, two-sided heat kernel estimates and two-sided resistance estimates are indeed equivalent. This equivalence was proved in [12] on graphs, in [45] on graphs and resistance forms and in [37] on metric spaces using an analytical approach.
@@
Lemma 8.4.
(Morrey-Sobolev inequality) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and PI($\Psi$). Then we have the following Morrey-Sobolev inequality MS($\Phi,\Psi$). There exists $C_{MS}\in(0,+\infty)$ such that for any $u\in\mathcal{F}$, for $m$-a.e. $x,y\in X$ with $x\neq y$, we have
@@
The proof is standard using telescopic technique, see also [12, Page 1654].
@@
$|u_{B_{n+1}}-u_{B_{n}}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n+1}}|u-u_{B_{n}}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|\mathrm{d}m\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|^{2}\mathrm{d}m\right)^{1/2}$
@@
$|u_{B(x,r)}-u_{B(y,r)}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})$
[…]
@@
The upper bound is obvious by MS($\Phi,\Psi$). We use the idea of an analytical proof of [37, Theorem 3.4] to give the lower bound. By the spectral calculus, it is easy to see that
@@
Definition 8.5.
([44, DEFINITION 3.1]) Let $X$ be a set. A pair $(\mathcal{E},\mathcal{G})$ is called a resistance form on $X$ if the following five conditions are satisfied.
@@
    Let $\sim$ be the equivalence relation on $\mathcal{G}$ given by $u,v\in\mathcal{G}$, $u\sim v$ if and only if $u-v$ is a constant function. Then $(\mathcal{G}/\sim,\mathcal{E})$ is a Hilbert space.
@@
Recall that for a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, its extended Dirichlet space is defined as the family of all $m$-measurable functions $u$ which is finite $m$-a.e. and there exists an $\mathcal{E}$-Cauchy sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ such that $u_{n}\to u$ $m$-a.e.. The sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ is called an approximation sequence of $u\in\mathcal{F}_{e}$. We can extend $\mathcal{E}$ to $\mathcal{F}_{e}$ and $\mathcal{F}=\mathcal{F}_{e}\cap L^{2}(X;m)$. See [26, Section 1.5] for more details. We can also define the harmonicity of a function in the extended Dirichlet space as follows. Let $D$ be a non-empty open subset of $X$. We say that $u\in\mathcal{F}_{e}$ is harmonic in $D$ if $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}\cap C_{c}(D)$.
@@
    Let $u\in\mathcal{F}_{e}$ be harmonic in a non-empty open subset $D$ of $X$. Then $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}_{e}\subseteq C(X)$ with $\varphi=0$ on $X\backslash D$.
@@
(2) Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent, by [26, Theorem 1.6.3], we have $1\in\mathcal{F}_{e}$ and $\mathcal{E}(1,1)=0$. For $u\in\mathcal{F}_{e}$, if $\mathcal{E}(u,u)=0$, then by MS($\Phi,\Psi$), we have $u$ is a constant function. Hence we have (RF1).
@@
By MS($\Phi,\Psi$) again, using the Arzelà-Ascoli lemma and the diagonal argument, there exists a subsequence, still denoted by $\{u_{n}\}$, and $u\in C(X)$ such that $u_{n}$ converges uniformly to $u$ on any relatively compact open subset of $X$. By the definition of $\mathcal{F}_{e}$, we have $u\in\mathcal{F}_{e}$ and $\{u_{n}\}$ is $\mathcal{E}$-convergent to $u$. Hence $(\mathcal{F}_{e}/\sim,\mathcal{E})$ is complete, hence we have (RF2).
@@
(RF5) follows from [26, Corollary 1.6.3].
@@
(4) Since $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$, by [44, PROPOSITION 9.13], we have the continuity and the quasi-continuity coincide. Then $\mathcal{F}_{e,D}$ given by [26, Equation (2.3.14)] can be re-written as
$\mathcal{F}_{e,D}:=\left\{u\in\mathcal{F}_{e}:\widetilde{u}=0\text{ q.e. on }X\backslash D\right\}=\left\{u\in\mathcal{F}_{e}\subseteq C(X):u=0\text{ on }X\backslash D\right\},$
@@
Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is regular, by [26, Exercise 1.4.1], we have $\mathcal{F}\cap C_{c}(X)$ is a special standard core of $\mathcal{E}$. Then (4) follows from the equivalence in [26, Theorem 2.3.3 (ii)]. ∎
By the resistance form theory, we can solve “boundary value problems” or take a trace of resistance forms as follows.
Lemma 8.7.
([44, Section 8]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $Y\subsetneqq X$ be a non-empty closed subset and $\mathcal{F}_{e}|_{Y}=\{u|_{Y}:u\in\mathcal{F}_{e}\}$. Then for any $u\in\mathcal{F}_{e}$, there exists a unique $v\in\mathcal{F}_{e}$ which is harmonic in $X\backslash Y$ such that $v=u$ on $Y$. Moreover, $v$ is the unique function in $\mathcal{F}_{e}$ that minimizes the following variational problem
@@
Corollary 8.8.
(Maximum principle) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $D\subsetneqq X$ be a non-empty open subset of $X$. Let $u,v\in\mathcal{F}_{e}$ be harmonic in $D$ and satisfy $u\leq v$ on $X\backslash D$, then $u\leq v$ in $D$. In particular, $\inf_{X\backslash D}u\leq u\leq\sup_{X\backslash D}u$.
@@
We only need to show that if $u\in\mathcal{F}_{e}$ is harmonic in $D$ and satisfies $u\leq 0$ on $X\backslash D$, then $u\leq 0$ in $D$. Indeed, $u\wedge 0\in\mathcal{F}_{e}$ satisfies $u\wedge 0=u$ on $X\backslash D$, by Lemma 8.7, we have
@@
The effective resistance between certain two sets can be achieved by the energy of some functions as follows.
@@
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $A,B$ be two non-empty closed subsets of $X$ satisfying $A\cap B=\emptyset$ and $A\cup B\subsetneqq X$. If one of the following conditions holds.
@@
Since $B$ is compact, $X\backslash A$ is open and $B\subseteq X\backslash A$, by the regularity of $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, there exists $u\in\mathcal{F}\cap C_{c}(X)\subseteq\mathcal{F}_{e}$ such that $0\leq u\leq 1$ in $X$, $u=1$ on $B$ and $\mathrm{supp}(u)\subseteq X\backslash A$. Hence $u|_{A\cup B}=1_{B}\in\mathcal{F}_{e}|_{A\cup B}$, $(1-u)|_{A\cup B}=1_{A}\in\mathcal{F}_{e}|_{A\cup B}$.
@@
where the RHS is equal to $R(A,B)^{-1}$ by definition. Since $\mathcal{F}\subseteq\mathcal{F}_{e}$, we only need to show $\geq$”. By the maximum principle Corollary 8.8, we have $0\leq v\leq 1$ in $X$.
For (a), since $0\leq v\leq 1$ in $X$, $v=0$ on $A$ and $X\backslash A$ is a relatively compact open set, we have $v\in\mathcal{F}_{e}\cap L^{2}(X;m)=\mathcal{F}$, $\geq$” is obvious.
For (b), since $A,B$ are both compact, by a similar proof to Lemma 8.6 (2), there exists an approximation sequence $\{v_{n}\}\subseteq\mathcal{F}\subseteq C(X)$ of $v\in\mathcal{F}_{e}$ that converges uniformly to $v$ on $A\cup B$. For any $\varepsilon\in(0,1/4)$, there exists $N\geq 1$ such that for any $n\geq N$, we have $|v_{n}-v|<\varepsilon$ on $A\cup B$, hence $-\varepsilon<v_{n}<\varepsilon$ on $A$ and $1-\varepsilon<v_{n}<1+\varepsilon$ on $B$. Let $\varphi:\mathbb{R}\to\mathbb{R}$ be given by
@@
that is, $\varphi\in C(\mathbb{R})$ satisfies $\varphi=0$ on $(-\infty,\varepsilon]$, $\varphi=1$ on $[1-\varepsilon,+\infty)$ and $\varphi$ is linear on $[\varepsilon,1-\varepsilon]$. Then $\varphi(0)=0$ and $|\varphi(t)-\varphi(s)|\leq\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}|t-s|$ for any $t,s\in\mathbb{R}$. For any $n\geq N$, let $\widetilde{v}_{n}=\varphi(v_{n})$, then $\widetilde{v}_{n}\in\mathcal{F}$ satisfies $\widetilde{v}_{n}=0$ on $A$, $\widetilde{v}_{n}=1$ on $B$ and $\mathcal{E}(\widetilde{v}_{n},\widetilde{v}_{n})\leq\left(\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\right)^{2}\mathcal{E}(v_{n},v_{n})$, hence
@@
We will be interested in the case $A=X\backslash B(x,r)$, $B=\{y\}$ for (a), and the case $A=\{x\}$, $B=\{y\}$ for (b), where $x,y\in X$, $r\in(0,+\infty)$.
@@
The proof is essentially the same as the proof of [12, LEMMA 2.4], [45, Lemma 4.1] and [37, Proposition 5.3].
@@
By V($\Phi$), there exists $N\geq 1$ depending only on $C_{VR},\alpha_{1},\alpha_{2},\varepsilon$, hence only on $C_{VR}$, $C_{R}$, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, such that $B\backslash\left(\frac{1}{2}B\right)$ can ben covered by at most $N$ balls with center in $B\backslash\left(\frac{1}{2}B\right)$ and radius $\varepsilon r$. Take $x_{1},\ldots,x_{N}\in B\backslash\left(\frac{1}{2}B\right)$ such that $B\backslash\left(\frac{1}{2}B\right)\subseteq\cup_{1\leq i\leq N}B(x_{i},\varepsilon r)$, let $h=\min_{1\leq i\leq N}h_{x_{i}}$. Then $h\in\mathcal{F}_{e}$ satisfies $h(x_{0})=1$, $0\leq h\leq 1$ in $X$, $h\leq\frac{1}{2}$ in $B\backslash\left(\frac{1}{2}B\right)$ and
@@
is the oscillation of the function $u\in l(A)$ on the set $A$.
The idea of the proof is from a recent result [40, Theorem 6.27], where a general result in the setting of $p$-energy forms was given to prove certain $p$-resistance defines indeed a metric.
@@
Let $Y=X\backslash B$, then $Y\subsetneqq X$ is a non-empty closed subset of $X$. For fixed $x,y\in B$, we have $u-u(x)\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, hence $u-u(x)=h_{Y\cup\{x\}}(u-u(x))$. By the maximum principle Corollary 8.8, we have $u-u(x)\leq(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, hence
@@
It remains to estimate $h_{Y\cup\{x\}}(1_{Y})(y)$. Indeed, by Corollary 8.9, we have
@@
Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{Y\cup\{x\}}(1_{\{x\}})=0$ on $Y\cup\{x\}$ and $h_{Y\cup\{x\}}(1_{Y})\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, by Lemma 8.6 (4), we have
@@
then $u_{1},v_{1}\in\mathcal{F}_{e}$ are harmonic in $B\backslash\{x,y\}$. Let $u_{2}=u_{1}|_{Y\cup\{x,y\}}$, $v_{2}=v_{1}|_{Y\cup\{x,y\}}$, then $u_{2},v_{2}\in\mathcal{F}_{e}|_{Y\cup\{x,y\}}$ and
@@
where $(\mathcal{E}|_{Y\cup\{x,y\}},\mathcal{F}_{e}|_{Y\cup\{x,y\}})$ is a resistance form on $Y\cup\{x,y\}$, which follows from Lemma 8.7. On $Y$, we have $u_{2}\in[0,1]$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}=u_{2}+v_{2}$. On $\{x\}$, we have $u_{2}=0$, $v_{2}=1$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=1=v_{2}=u_{2}+v_{2}$. On $\{y\}$, we have $u_{2}=0$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=0=u_{2}+v_{2}$. Hence $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}+v_{2}$. By the following easy result Lemma 8.13, we have

### v3

@@ 4 Poisson equation on unbounded MMD spaces
[…]
@@
In this section, the strongly recurrent condition (1.3) is intrinsically required to ensure that the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ indeed induces a resistance form, as will be shown in Lemma 8.6, after which the arguments rely heavily on Kigami’s resistance form theory [46].
Recall that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent if for any non-negative function $f\in L^{1}(X;m)$, we have either $Gf=+\infty$ or $Gf=0$ mm-a.e., where $G$ is the Green operator, see [27, Equation (1.6.2)]. In our setting, the Dirichlet form is recurrent as follows.
@@
To begin with, let us introduce effective resistances. Let $A,B$ be two non-empty subsets of $X$. The effective resistance between $A$ and $B$ is defined as follows.
@@
In the strongly recurrent setting, two-sided heat kernel estimates and two-sided resistance estimates are indeed equivalent. This equivalence was proved in [13] for graphs, in [47] for graphs and resistance forms, and in [39] for metric spaces using an analytical approach.
@@
Lemma 8.4 (Morrey-Sobolev inequality).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and PI($\Psi$). Then we have the following Morrey-Sobolev inequality MS($\Phi,\Psi$). There exists $C_{MS}\in(0,+\infty)$ such that for any $u\in\mathcal{F}$, for $m$-a.e. $x,y\in X$ with $x\neq y$, we have
@@
The proof is standard using a telescopic technique, see also [13, Page 1654].
@@
$|u_{B_{n+1}}-u_{B_{n}}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n+1}}|u-u_{B_{n}}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n}}|u-u_{B_{n}}|\mathrm{d}m\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n}}|u-u_{B_{n}}|^{2}\mathrm{d}m\right)^{1/2}$
@@
$|u_{B(x,r)}-u_{B(y,r)}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})$
$\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|^{2}m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})\right)^{1/2}$
@@
The upper bound is obvious by MS($\Phi,\Psi$). We use the idea of an analytical proof of [39, Theorem 3.4] to give the lower bound. By the spectral calculus, it is easy to see that
@@
Definition 8.5 ([46, DEFINITION 3.1]).
Let $X$ be a set. A pair $(\mathcal{E},\mathcal{G})$ is called a resistance form on $X$ if the following five conditions are satisfied.
@@
    Let $\sim$ be the equivalence relation on $\mathcal{G}$ given by $\in\mathcal{G}, uvu\sim v$ if and only if uvu-v is a constant function. Then $(\mathcal{G}/\!\sim,\mathcal{E})$ is a Hilbert space.
@@
Recall that for a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, its extended Dirichlet space is defined as the family of all $m$-measurable functions $u$ which is finite $m$-a.e. and there exists an $\mathcal{E}$-Cauchy sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ such that $u_{n}\to u$ $m$-a.e.. The sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ is called an approximation sequence of $u\in\mathcal{F}_{e}$. We can extend $\mathcal{E}$ to $\mathcal{F}_{e}$ and $\mathcal{F}=\mathcal{F}_{e}\cap L^{2}(X;m)$. See [27, Section 1.5] for more details. We can also define the harmonicity of a function in the extended Dirichlet space as follows. Let $D$ be a non-empty open subset of $X$. We say that $u\in\mathcal{F}_{e}$ is harmonic in $D$ if $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}\cap C_{c}(D)$.
@@
    Let $u\in\mathcal{F}_{e}$ be harmonic in a non-empty open subset $D$ of $X$. Then $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}_{e,D}$, where
    $\mathcal{F}_{e,D}=\left\{u\in\mathcal{F}_{e}\subseteq C(X):u=0\text{ on }X\backslash D\right\}.$ (8.3)
@@
(2) Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent, by [27, Theorem 1.6.3], we have $1\in\mathcal{F}_{e}$ and $\mathcal{E}(1,1)=0$. For $u\in\mathcal{F}_{e}$, if $\mathcal{E}(u,u)=0$, then by MS($\Phi,\Psi$), we have $u$ is a constant function. Hence we have (RF1).
@@
By MS($\Phi,\Psi$) again, using the Arzelà-Ascoli lemma and the diagonal argument, there exists a subsequence, still denoted by $\{u_{n}\}$, and $u\in C(X)$ such that $u_{n}$ converges uniformly to $u$ on any relatively compact open subset of $X$. By the definition of $\mathcal{F}_{e}$, we have $u\in\mathcal{F}_{e}$ and $\{u_{n}\}$ is $\mathcal{E}$-convergent to $u$. Hence $(\mathcal{F}_{e}/\!\sim,\mathcal{E})$ is complete, then we have (RF2).
@@
(RF4) follows from MS($\Phi,\Psi$).
(RF5) follows from [27, Corollary 1.6.3].
@@
(4) Since $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$, by [46, PROPOSITION 9.13], we have the continuity and the quasi-continuity coincide. By [27, Equation (2.3.14)], we have
$\mathcal{F}_{e,D}=\left\{u\in\mathcal{F}_{e}:\widetilde{{u}}=0\text{ q.e. on }X\backslash D\right\},$
@@
Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is regular, by [27, Exercise 1.4.1], we have $\mathcal{F}\cap C_{c}(X)$ is a special standard core of $\mathcal{E}$. Then (4) follows from the equivalence in [27, Theorem 2.3.3 (ii)]. ∎
By resistance form theory, we can solve “boundary value problems” and take traces of resistance forms as follows.
Lemma 8.7 ([46, Section 8]).
[…]
@@
Corollary 8.8 (Maximum principle).
[…]
@@
[…]
@@
For certain pairs of sets, the effective resistance can be realized by the energy of a suitable function, as follows.
@@
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $A,B$ be two non-empty closed subsets of $X$ satisfying $A\cap B=\emptyset$ and $A\cup B\subsetneqq X$. Assume that one of the following conditions holds:
@@
[…]
@@
where the RHS is equal to $R(A,B)^{-1}$ by definition. Since $\mathcal{F}\subseteq\mathcal{F}_{e}$, we only need to show $\geq$”. By the maximum principle (Corollary 8.8), we have $0\leq v\leq 1$ in $X$.
[…]
Assuming (b), since $A,B$ are both compact, by a similar proof to Lemma 8.6 (2), there exists an approximation sequence $\{v_{n}\}\subseteq\mathcal{F}\subseteq C(X)$ of $v\in\mathcal{F}_{e}$ that converges uniformly to $v$ on $A\cup B$. For any $\varepsilon\in(0,1/4)$, there exists $N\geq 1$ such that for any $n\geq N$, we have $|v_{n}-v|<\varepsilon$ on $A\cup B$, hence $\varepsilon<v_{n}<\varepsilon$ on $A$ and $1-\varepsilon<v_{n}<1+\varepsilon$ on $B$. Let $\varphi:\mathbb{R}\to\mathbb{R}$ be given by
@@
that is, $\varphi\in C(\mathbb{R})$ satisfies $\varphi=0$ on $(-\infty,\varepsilon]$, $\varphi=1$ on $1-\varepsilon,+\infty)$ and $\varphi$ is linear on $\varepsilon,1-\varepsilon]$. Then $\varphi(0)=0$ and $|\varphi(t)-\varphi(s)|\leq\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}|t-s|$ for any $t,s\in\mathbb{R}$. For any $n\geq N$, let $\widetilde{{v}}_{n}=\varphi(v_{n})$, then $\widetilde{{v}}_{n}\in\mathcal{F}$ satisfies $\widetilde{{v}}_{n}=0$ on $A$, $\widetilde{{v}}_{n}=1$ on $B$ and $\mathcal{E}(\widetilde{{v}}_{n},\widetilde{{v}}_{n})\leq\left(\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\right)^{2}\mathcal{E}(v_{n},v_{n})$, hence
@@
[…]
@@
The proof is essentially the same as the proof of [13, LEMMA 2.4], [47, Lemma 4.1] and [39, Proposition 5.3].
@@
By V($\Phi$), there exists $N\geq 1$ depending only on $C_{VR},\alpha_{1},\alpha_{2},\varepsilon$, hence only on $C_{VR}$, $C_{R}$, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, such that $B\backslash\left(\frac{1}{2}B\right)$ can be covered by at most $N$ balls with center in $B\backslash\left(\frac{1}{2}B\right)$ and radius $\varepsilon r$. Take $x_{1},\ldots,x_{N}\in B\backslash\left(\frac{1}{2}B\right)$ such that $B\backslash\left(\frac{1}{2}B\right)\subseteq\cup_{1\leq i\leq N}B(x_{i},\varepsilon r)$, let $h=\min_{1\leq i\leq N}h_{x_{i}}$. Then $h\in\mathcal{F}_{e}$ satisfies $h(x_{0})=1$, $0\leq h\leq 1$ in $X$, $h\leq\frac{1}{2}$ in $B\backslash\left(\frac{1}{2}B\right)$ and
@@
is the oscillation of the function $u\in l(A)$ on the set AA.
The idea of the following proof is from a recent result [42, Theorem 6.27], where a general result in the setting of $p$-energy forms was given to prove certain $p$-resistance defines indeed a metric. An alternative proof suggested by Kajino and Shimizu will be provided afterward.
@@
[…]
@@
$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\geq-h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y}).$
It suffices to estimate $h_{Y\cup\{x\}}(1_{Y})(y)$. Indeed, by Corollary 8.9, we have
@@
[…]
@@
[…]
@@
where $(\mathcal{E}|_{Y\cup\{x,y\}},\mathcal{F}_{e}|_{Y\cup\{x,y\}})$ is a resistance form on $Y\cup\{x,y\}$, which follows from Lemma 8.7. On $Y$, we have $u_{2}\in[0,1]$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}=u_{2}+v_{2}$. On $\{x\}$, we have $u_{2}=0$, $v_{2}=1$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=1=v_{2}=u_{2}+v_{2}$. On $\{y\}$, we have $u_{2}=0$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=0=u_{2}+v_{2}$. Hence $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}+v_{2}$. By the following easy result, Lemma 8.13, we have